OpenStudy (anonymous):

simplify under root 7(underroot 42 -underroot 98)/7

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt7 (\sqrt{42}-\sqrt{98})}{7}\] this ? :)

Parth (parthkohli):

I know how to solve this.

OpenStudy (anonymous):

I am asking... dear asker, is this what you mean by your "word" problem? (my first post) ?

OpenStudy (anonymous):

@shafaq

OpenStudy (radar):

this is:\[\sqrt{6}-\sqrt{2}\] Does this equal\[\sqrt{7}(\sqrt{42}-\sqrt{98})\over7\] being from Missouri "show me"

OpenStudy (anonymous):

its \[\sqrt{2} (\sqrt{3} - \sqrt{7})\]

OpenStudy (radar):

\[\sqrt{7}(\sqrt{7}\sqrt{6}-\sqrt{7}\sqrt{7}\sqrt{2})\over7\] \[7\sqrt{6}-7\sqrt{7}\sqrt{2}\over7\] \[\sqrt{6}-\sqrt{7}\sqrt{2}\] \[\sqrt{3}\sqrt{2}-\sqrt{7}\sqrt{2}\]\[\sqrt{2}(\sqrt{3}-\sqrt{7})\]In case you wanted to see how we got there.

OpenStudy (radar):

vedic has a good answer

OpenStudy (anonymous):

@radar thnku..:)

OpenStudy (radar):

@ParthKohli, could you show step by step how you reached that solution

Parth (parthkohli):

MY ANSWER IS CORRECT http://www.wolframalpha.com/input/?i=sqrt6+-+sqrt2

Parth (parthkohli):

\[\LARGE \sqrt{2}(\sqrt{3} - 1)\]

Parth (parthkohli):

Can I have a medal now?

OpenStudy (anonymous):

@ParthKohli its incorrect...

OpenStudy (radar):

I don't see how \[\sqrt{2}(\sqrt{3}-1)\]could be a solution, I see that it has been deleted, was there an error found?

Parth (parthkohli):

sqrt6 - sqrt2 = sqrt2 * sqrt3 - sqrt2 Factorizing, sqrt2(sqrt3 - 1) I deleted it because I thought that I was wrong because the majority had the same answer. I didn't wanna give a wrong idea to the asker. So, I deleted it.

OpenStudy (radar):

I was wondering where it strayed. I made multiple errors in my attempt to solve. It was a tricky procedure.

Parth (parthkohli):

May I have a medal now?

OpenStudy (radar):

I can award only one medal. lol The new rules and software only allow one award per question.

Parth (parthkohli):

Lol then undo that medal and give it to me..

OpenStudy (radar):

They may change it back, it is not too popular.

OpenStudy (anonymous):

how parth answer is correct xplain me...

OpenStudy (radar):

Sorry but vedic had it right, still would of liked to seen a step by step progression that resulted in the solution.

Parth (parthkohli):

I did explain here, see all of my answers

OpenStudy (radar):

Your factorization was o.k. but the final results were wrong.

Parth (parthkohli):

sqrt6 - sqrt2 = sqrt2 * sqrt3 - sqrt2 Factorizing, sqrt2(sqrt3 - 1) Here it is again

OpenStudy (anonymous):

how u arrived at sqrt6 - sqrt2.... expalin..

OpenStudy (radar):

That was your solution, mine was\[\sqrt{2}(\sqrt{3}-\sqrt{7})\]

OpenStudy (anonymous):

hellooo... Can I know what is the given please? O_O :(

Parth (parthkohli):

Just like everyone did.. sqrt7(sqrt 42 - sqrt98) all over 7 7sqrt6 - 7sqrt2 all over 7 7(sqrt6 - sqrt2) all over 7 [both 7's would be cut] sqrt6 - sqrt2 THAT's how I arrived. MEDAL?

OpenStudy (anonymous):

@ParthKohli sqrt7 must be multiplied by both..

Parth (parthkohli):

lol that's what I did in 2nd step

OpenStudy (anonymous):

\[\sqrt{98} =\sqrt{7\times7\times2}\]

OpenStudy (radar):

Don't you see the error?

OpenStudy (anonymous):

so one sqrt 7 from outside makes it 3*sqrt7

OpenStudy (radar):

@Kreshnik, the question is posted above by the asker. There has been two answers presented. Although, one was deleted.

OpenStudy (anonymous):

but I DO NOT understand the question posted by the asker.... is it like my first post? :(

Parth (parthkohli):

Yes @Kreshnik

OpenStudy (radar):

Yes your first post seems to be correct, and it is what I used to solve it as it was very clear.

OpenStudy (anonymous):

@ParthKohli u found ur mistake or still believe me and radar are wrong?

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt{7}(\sqrt{42}-\sqrt{98})}{7}=\] \[\LARGE \frac{(\sqrt{42\cdot 7}-\sqrt{98\cdot 7 })}{7}=\] \[\LARGE \frac{(\sqrt{49\cdot 6}-\sqrt{49\cdot 14 })}{7}=\] \[\LARGE \frac{(7\sqrt{ 6}-7\sqrt{ 14 })}{7}=\] \[\LARGE \frac{\cancel7(\sqrt{ 6}-\sqrt{ 14 })}{\cancel 7}=...\]

OpenStudy (anonymous):

@radar seems to be right :)

OpenStudy (radar):

Thank you for your support.

OpenStudy (radar):

@Kreshnik, I like that method as it leaves less chance for errors.

OpenStudy (radar):

Plus, it is easy to read, the fonts are nice.

OpenStudy (radar):

@shafaq, I believe you question has been answered.